OpenStudy (anonymous):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (3, 6) and (1, −2)?

OpenStudy (anonymous):

hint : the function of a line is f(x) = ax + b

OpenStudy (anonymous):

It's not a function though, it's a slope.

OpenStudy (anonymous):

Find the gradient of the points given?

OpenStudy (anonymous):

It's trying to find the slope.

OpenStudy (anonymous):

gradient is also called as the slope..so find the gradient of the two points given

OpenStudy (anonymous):

I don't know how to do that, that's why I have it on here. It was on a test, and it was one of the questions I got wrong. So I just wanna know how to do it right.

OpenStudy (anonymous):

6 = 3a + b -2 = a + b

OpenStudy (anonymous):

okay the gradient of a line, when two points are given can be found using the following equation \[(x_{1}, y_{1}) and (x_{2},y_{2})\] \[m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (anonymous):

find a and b... a is the slope

OpenStudy (anonymous):

\[m = \frac{ -2 - 6 }{ 1-3 }\] Is that right @thushananth01

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

do I simplify?

OpenStudy (anonymous):

you hve to :)

OpenStudy (anonymous):

so then when I simplify, I got -4/-1, wouldn't that just be -4?

OpenStudy (anonymous):

@thushananth01

OpenStudy (anonymous):

No it's 4

OpenStudy (anonymous):

- / - = +

OpenStudy (anonymous):

ooooooh okay. Thanks :)